# KCL-PH-TH/2024-38

### v2

@@ 3.1 Model and particle production mechanism
×$\times{\rm csch}^{2}\left(\frac{\pi L_{w}\Delta p^{z}}{2}\right)\,,$ (20)
@@
×$\times\int_{-|k^{z}|_{\rm max}(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|)}^{|k^{z}|_{\rm max}(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|)}\mathrm{d}k^{z}\,\frac{1}{E_{\mathbf{k}}^{(\chi)}}\frac{1}{\sqrt{G(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|,k^{z})}}{\rm csch}^{2}\left(\frac{\pi L_{w}\overline{\Delta p^{z}}}{2}\right)\,,$ (24)
@@
×$\times\int_{-|k^{z}|_{\rm max}(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|)}^{|k^{z}|_{\rm max}(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|)}\mathrm{d}k^{z}\,\frac{1}{E_{\mathbf{k}}^{(\chi)}}\frac{1}{\sqrt{G(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|,k^{z})}}{\rm csch}^{2}\left(\frac{\pi L_{w}\overline{\Delta p^{z}}}{2}\right)\,,$ (28)
@@ 4.3 Numerical results and fit
$\boxed{n_{V}\approx 6.9\times 10^{-7}\times\frac{\gamma_{w}T^{4}\lambda^{2}v_{b}^{2}}{L_{w}m_{V}^{4}}\,.}$ (60)
@@
$\boxed{\mkern 1.5mu\overline{\mkern-1.5mu\Omega\mkern-1.5mu}\mkern 1.5mu_{V,\rm BE}^{\rm today}h^{2}\approx 4.1\times\left(\frac{\gamma_{w}}{10^{7}}\right)\left(\frac{\lambda}{0.1}\right)^{2}\left(\frac{10}{L_{w}T_{n}}\right)\left(\frac{100T_{n}}{m_{V}}\right)^{2}\left(\frac{100v_{b}}{m_{V}}\right)\left(\frac{100}{g_{\star S}(T_{\rm after})}\right)\left(\frac{v_{b}}{100\rm GeV}\right)\left(\frac{T_{n}}{T_{\rm after}}\right)^{3}\,.}$ (61)
@@ 5.2 Friction from vector dark matter particle production $\mathcal{P}_{\phi\rightarrow 2V}$
One may compare this result with that given in Ref. [85]. In Eq. (72) of Ref. [85], noting that $n_{h}\sim T^{3}$ and that Ref. [85] assumes $L_{w}\sim 1/v_{b}$, one gets the same dependence on $v_{b}$, $\gamma_{w}$, $T$ and $L_{w}$. The superscript “$(2)$” reminds us that this expression is only valid in the quadratic scaling regime defined by Eq. (70). For larger $\gamma_{w}$, the quadratic scaling may be replaced by the linear-logarithmic scaling, Eq. (69). However, we note that the scanned range for $\gamma_{w}$ in the above fitting procedure does not go beyond the quadratic scaling regime; the quadratic scaling may therefore be valid for higher $\gamma_{w}$ and could even be exact. [^7] The reason we chose the logarithmic function is that the same logarithmic structure for the scalar case, $\mathcal{P}_{\phi\rightarrow 2\chi}$, found in Ref. [82] has analytic support [14]. We leave this open question for future study. We also note that a quadratic behaviour, valid up to $\gamma_{w}\sim 1/(m_{V}L_{w})$, was previously observed for the friction $\mathcal{P}_{V\rightarrow V}$ caused by the 1-to-1 process of the vector boson crossing the wall, which is associated with a factor of $\rho_{V}(\Delta m^{2}_{V}/m^{2}_{V})^{2}$ where $\rho_{V}$ is the energy density of the vector boson outside of the bubble [115]. In our case, this contribution is negligible due to the suppression from both $\rho_{V}$ and $(\Delta m^{2}_{V}/m^{2}_{V})^{2}$ for large $m_{V}$.
@@ 5.3 Terminal wall velocity in the ballistic-detonation regime
[…]
@@ 6 Parameter space of heavy WIMP vector dark matter
Having shown that ultra-relativistic bubble expansion during a FOPT is both possible and can efficiently source a number density of heavy vector dark matter, we now consider the implications of this new production mechanism on the viability of TeV-scale WIMP vector dark matter. Specifically, we focus on the region of parameter space where thermal freeze-out is unable to generate the observed relic abundance. For non-thermal production mechanisms of vector dark matter to dominate over thermal freeze-out, we require its scalar coupling to be sufficiently large for the thermal relic density to be negligible. Non-thermal bubble wall production at the time of the phase transition could then generate a greater abundance than the observed relic density. However, its subsequent Boltzmann evolution must be taken into account which can reduce the non-thermally produced dark matter to its observed relic abundance. Planck constraints on $\Omega_{\rm{M}0}h^{2}$ and $\Omega_{{\rm B}0}h^{2}$ tell us that $\Omega_{{\rm DM}0}h^{2}=0.121$ [119] is a good approximation to a couple of percent, which is a good enough approximation for this work. The subscripts “M”, “DM” and “B” correspond to matter, dark matter and baryons, respectively, and as usual the subscript 00 refers to quantities measured at $t=t_{0}$, (i.e. today, at redshift $z=0$). The dark matter relic abundance is

### v3

@@ 3.1 Model and particle production mechanism
$\times{\rm csch}^{2}\left(\frac{\pi L_{w}\Delta p^{z}}{2}\right)\,$, (20)
@@
$\times\int_{-|k^{z}|_{\rm max}(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|)}^{|k^{z}|_{\rm max}(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|)}\mathrm{d}k^{z}\,\frac{1}{E_{\mathbf{k}}^{(\chi)}}\frac{1}{\sqrt{G(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|,k^{z})}}{\rm csch}^{2}\left(\frac{\pi L_{w}\overline{\Delta p^{z}}}{2}\right)\,$, (24)
@@
$\times\int_{-|k^{z}|_{\rm max}(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|)}^{|k^{z}|_{\rm max}(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|)}\mathrm{d}k^{z}\,\frac{1}{E_{\mathbf{k}}^{(\chi)}}\frac{1}{\sqrt{G(|\mathbf{p}_{\perp}|,|\mathbf{k}_{\perp}|,k^{z})}}{\rm csch}^{2}\left(\frac{\pi L_{w}\overline{\Delta p^{z}}}{2}\right)\,$, (28)
@@ 4.3 Numerical results and fit
$\boxed{n_{V}\approx 6.9\times 10^{-4}\times\frac{\gamma_{w}T^{4}\lambda^{2}v_{b}^{2}}{L_{w}m_{V}^{4}}\,.}$ (60)
@@
$\boxed{\mkern 1.5mu\overline{\mkern-1.5mu\Omega\mkern-1.5mu}\mkern 1.5mu_{V,\rm BE}^{\rm today}h^{2}\approx 4.1\times\left(\frac{\gamma_{w}}{10^{4}}\right)\left(\frac{\lambda}{0.1}\right)^{2}\left(\frac{10}{L_{w}T_{n}}\right)\left(\frac{100T_{n}}{m_{V}}\right)^{2}\left(\frac{100v_{b}}{m_{V}}\right)\left(\frac{100}{g_{\star S}(T_{\rm after})}\right)\left(\frac{v_{b}}{100\rm GeV}\right)\left(\frac{T_{n}}{T_{\rm after}}\right)^{3}\,.}$ (61)
@@ 5.2 Friction from vector dark matter particle production $\mathcal{P}_{\phi\rightarrow 2V}$
One may compare this result with that given in Ref. [18]. In Eq. (72) of Ref. [18], noting that $n_{h}\sim T^{3}$ and that Ref. [18] assumes $L_{w}\sim 1/v_{b}$, one gets the same dependence on $v_{b}$, $\gamma_{w}$, $T$ and $L_{w}$. The superscript “(2)(2)” reminds us that this expression is only valid in the quadratic scaling regime defined by Eq. (70). For larger $\gamma_{w}$, the quadratic scaling may be replaced by the linear-logarithmic scaling, Eq. (69). However, we note that the scanned range for $\gamma_{w}$ in the above fitting procedure does not go beyond the quadratic scaling regime; the quadratic scaling may therefore be valid for higher $\gamma_{w}$ and could even be exact. [^7] The reason we chose the logarithmic function is that the same logarithmic structure for the scalar case, $\mathcal{P}_{\phi\rightarrow 2\chi}$, found in Ref. [13] has analytic support [29]. We leave this open question for future study. We also note that a quadratic behaviour, valid up to $\gamma_{w}\sim 1/(m_{V}L_{w})$, was previously observed for the friction $\mathcal{P}_{V\rightarrow V}$ caused by the 1-to-1 process of the vector boson crossing the wall, which is associated with a factor of $\rho_{V}(\Delta m^{2}_{V}/m^{2}_{V})^{2}$ where $\rho_{V}$ is the energy density of the vector boson outside of the bubble [63]. In our case, this contribution is negligible due to the suppression from both $\rho_{V}$ and $(\Delta m^{2}_{V}/m^{2}_{V})^{2}$ for large $m_{V}$.
@@ 5.3 Terminal wall velocity in the ballistic-detonation regime
Transition radiation friction, with its linear dependence in $\gamma_{w}$, severely restricts the terminal wall velocity. Even if we assume that the wall velocity gets past the hydrodynamic obstruction and enters into the ballistic-detonation regime, a large Lorentz factor $\gamma_{w}$ requires a very strong supercooling in the presence of transition radiation. According to Ref. [68], for a first-order electroweak phase transition, one has $\gamma_{w}\sim\alpha_{n}^{3/4}$. [^8] See Eq. (6.10) of Ref. [68] and notice that $T_{\rm start}/T_{\rm nuc}\approx\alpha_{n}^{1/4}$. On the other hand, there is a maximal phase transition strength $\alpha_{n}$ above which the phase transition can never complete [75, 56]. This makes large Lorentz factors unlikely for a first-order electroweak phase transition. See also Ref. [32] for an analysis based on the Standard Model effective field theory where it is shown that only $\gamma_{w}\;\raisebox{-3.00003pt}{$\stackrel{{\scriptstyle\displaystyle<}}{{\sim}}$}\;10$ is feasible. However, if we consider a dark sector FOPT where $\phi$ couples only with the dark matter (and perhaps also the Higgs field), then the transition-radiation process is absent and it is possible to have very large Lorentz factors $\gamma_{w}$. We now show that this is still the case when including friction from vector boson pair production, despite a quadratic scaling in $\gamma_{w}$.
@@ 6 Parameter space of heavy WIMP vector dark matter
Having shown that ultra-relativistic bubble expansion during a FOPT is both possible and can efficiently source a number density of heavy vector dark matter, we now consider the implications of this new production mechanism on the viability of TeV-scale WIMP vector dark matter. Specifically, we focus on the region of parameter space where thermal freeze-out is unable to generate the observed relic abundance. For non-thermal production mechanisms of vector dark matter to dominate over thermal freeze-out, we require its scalar coupling to be sufficiently large for the thermal relic density to be negligible. Non-thermal bubble wall production at the time of the phase transition could then generate a greater abundance than the observed relic density. However, its subsequent Boltzmann evolution must be taken into account which can reduce the non-thermally produced dark matter to its observed relic abundance. Planck constraints on $\Omega_{\rm{M}0}h^{2}$ and $\Omega_{{\rm B}0}h^{2}$ tell us that $\Omega_{{\rm DM}0}h^{2}=0.121$ [3] is a good approximation to a couple of percent, which is a good enough approximation for this work. The subscripts “M”, “DM” and “B” correspond to matter, dark matter and baryons, respectively, and as usual the subscript 0 refers to quantities measured at $t=t_{0}$, (i.e. today, at redshift $z=0$). The dark matter relic abundance is